OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)

OpenStudy (anonymous):

This is simple. Use your slope formula \[y1-y2/ x1-x2\]

OpenStudy (anonymous):

The slopes are parrallel, meaning the slopes are the same.

OpenStudy (anonymous):

can u help me put it together please

OpenStudy (anonymous):

(7--8)/ (3--2) = (7+8)/(3+2) = 15/5 = 3. Which is your slope. The other coordinate can be used to find your equation of your line.

OpenStudy (anonymous):

so what now

OpenStudy (anonymous):

3 is the slope

OpenStudy (anonymous):

The question asked for slope correct? 3 is your slope since the lines are parrelell meaning they have the same slope. Yes

OpenStudy (anonymous):

i gotcha :)